# Flexible and efficient spatial extremes emulation via variational autoencoders

### v3

@@ 3 Simulation study
In this section, we simulate data from five different parametric models that have varying levels of extremal dependence across space. By examining the diagnostics introduced in Section 2, we validate the efficacy of our XVAE to analyze and emulate data from both model (3) and misspecified models.
@@ 3.1 General setting
To assess the performance of our XVAE, we conduct a simulation study in which data are generated at $n_{s}=2,000$ random locations uniformly sampled over the square $[0,10]\times[0,10]$. We simulate $n_{t}=100$ replicates of the process from each of the following different models:
@@
When simulating from Models IIIV, we use time-invariant dependence parameters $\alpha_{t}\equiv\alpha$ and $\boldsymbol{\gamma}_{t}\equiv\boldsymbol{\gamma}$; see Figure 1 for the knot locations and $\boldsymbol{\gamma}$ values. Recall that $K$ is the number of basis functions and $\mathcal{D}=\{k:\gamma_{k}=0\}$. Models IV gradually exhibit increasingly stronger extremal dependence, and they can help us test whether the XVAE can capture spatially-varying dependence structures that exhibit local AD and/or local AI. Since the proposed process (3) allows $\gamma_{k}$ to change at different knots ($k=1,\ldots,K$), a well-trained XVAE should be able to differentiate between local AD ($\gamma_{k}=0$) and local AI ($\gamma_{k}>0$).
Figure 1: The left panel presents knot locations used for simulating data under Models IIIV, and we only show the support of the one Wendland basis function centered at knot in the middle of the domain. Model V uses the same set of knots but the basis functions are not compactly supported. The middle and right panels display the $\gamma_{k}$ values, $k=1,\ldots,K$, used in the exponentially-tilted PS variables at each knot for Models II and III respectively. The circled knots signify $\gamma_{k}=0$, which induces local AD.
Model I is a stationary and isotropic Gaussian process with a Matérn covariance function. It is known that the joint distribution of the Gaussian process at any two locations $\boldsymbol{s}_{i}$ and $\boldsymbol{s}_{j}$ is light-tailed and leads to AI unless the correlation equals one. For Models II, III and IV, we simulate data from the max-id model (3) with $K=25$ evenly-spread knots across the grid, denoted by $\{\tilde{\boldsymbol{s}}_{1},\ldots,\tilde{\boldsymbol{s}}_{K}\}$. Setting the range parameter to $r=3$, we use compactly supported Wendland basis functions $\omega_{k}(\boldsymbol{s},r)\propto\{1-d(\boldsymbol{s},\tilde{\boldsymbol{s}}_{k})/r\}^{2}_{+}$ centered at each knot (Wendland, 1995), $k=1,\ldots,K$; see Figure 1. The basis function values are standardized so that for each $\boldsymbol{s}$, $\sum_{k=1}^{K}\omega_{k}(\boldsymbol{s},r)=1$. The main difference between Models II, III and IV lies in the $\gamma_{k}$ values: Model II has no zero $\gamma_{k}$’s (i.e., $|\mathcal{D}|=0$), whereas Model III has a mix of positive and zero $\gamma_{k}$’s, and Model IV has only zero $\gamma_{k}$’s (i.e., $|\mathcal{D}|=K$). By Theorem 1.3, we know Model II gives only local AI and Model IV gives only local AD. In contrast, Model III gives both local AD and local AI. By contrast, Model V adopts the same set of knots but it uses Gaussian radial basis functions which are not compactly supported. Therefore, Model V is the Reich and Shaby (2012) max-stable model, and has stronger extremal dependence than Models IIV.
@@
When simulating from Models IIV, we choose the parameter $\alpha=1/2$ and then sample the latent variables $\boldsymbol{Z}_{t}$ from the exponentially-tilted PS distribution for each time replicate. Additionally, for Models IIIV, the white noise process $\epsilon_{t}(\boldsymbol{s})$ follows the same independent Fréchet($0,\tau,1/\alpha_{0}$) distribution with $\tau=1$ and $\alpha_{0}=1/4$.
@@
For each space-time simulated dataset, we randomly set aside 100 locations as a validation set. Subsequently, we analyze the dependency structure of the remaining 1,900 locations using both the proposed XVAE (initialized with data-driven knots unless specified otherwise) and a Gaussian process regression with heteroskedastic noise implemented in the R package hetGP (Binois and Gramacy, 2021). We then perform predictions at the 100 holdout locations as outlined in Section 2.1. In the following result sections, we show that both emulators will have good performances when emulating datasets from Models I and II, but only XVAE appropriately captures AD in Models IIIV.
@@ 3.2 Emulation results
Figure 2 and Figure E.1 of the Supplementary Material compare emulated replicates from XVAE and hetGP with data replicates from Models IV, while Figure E.2 displays QQ-plots that align well with the 1-1 line in all cases for XVAE but not for hetGP. Since the Gaussian process has much weaker extremal dependence, the resulting $\boldsymbol{\gamma}_{t}$ estimated in (8) after convergence is consistently far greater than $0.1$, indicating light tails in the exponentially-tilted PS variables and thus, local AI at all knots. Similarly, for Model II, there is AI everywhere in the domain. However, the $\boldsymbol{\gamma}_{t}$ values we used for Model II are much smaller than 0.1 (see Figure 1) and thus the fitted exponentially-tilted PS variables are heavier-tailed than the ones from Model I. Therefore, hetGP has difficulty in capturing the extremal dependence and the QQ-plot shows that large values in the tail tend to be underestimated, even though Model II still exhibits AI only.
@@
Figure 3 compares the performance of spatial predictions at the 100 holdout locations. For Model I, hetGP has lower CRPS and MSPE scores, indicating higher predictive power, as expected since the true process is Gaussian. However, the XVAE model still performs quite well in this case. For Models IIV, XVAE uniformly outperforms hetGP. Also, the CRPS and MSPE for hetGP are significantly higher for time replicates with extreme events.
The left three panels of Figure 4 and Figure E.3 of the Supplementary Material compare nonparametric estimates of the upper tail dependence $\chi_{h}(u)$ from the data replicates and emulations at three different distances $h\in\{0.5,2,5\}$ under the working assumption of stationarity. In general, we see that the dependence strength decays as $h$ and $u$ increase, with varying levels of positive limits as $u\rightarrow 1$ for Models IIIV. The results in Figure 4 demonstrate that our XVAE manages to accurately emulate the dependence behavior at both low and high quantiles and the empirical confidence envelopes of $\chi_{h}(u)$ are essentially indistinguishable between the simulated and emulated data.
@@
Figure 5: Initializing the XVAE using the true knots from Models II (top) and III (bottom), we show the estimates of $(\gamma_{1t},\gamma_{3t})^{T}$ (left) and $(\gamma_{5t},\gamma_{12t})^{T}$ (middle) from 1,000 samples generated with the trained decoder $(t=1)$. On the right, we also show the medians, 2.5% and 97.5% quantiles of the $n_{t}$ estimates of $\{\gamma_{kt}:k=1,\ldots,K\}$ for $t=1$, from the decoder (8), in which the 1-1 line is displayed in black for reference.
To showcase the inferential capabilities of our approach, we initialize the XVAE with true knots and rerun it on datasets simulated from Models II and III. Figure 5 displays $\boldsymbol{\gamma}_{t}$ estimates obtained by running the decoder (i.e., Eq. (8)) $1000$ times at $t=1$. The results highlight the XVAE’s ability to produce accurate estimates of $\boldsymbol{\gamma}_{t}=\{\gamma_{kt}:k=1,\ldots,K\}$, and correctly identify the extremal dependence class, with satisfactory agreement between true and estimated values, accounting for uncertainty. Additionally, we perform a coverage analysis by simulating $99$ more datasets with $n_{s}=2000$ and $n_{t}=100$ from Models II and III, running the XVAE on each to generate empirical credible intervals for $\boldsymbol{\gamma}_{t}$. Figure E.5 of the Supplementary Material shows the coverage probabilities of $\{\gamma_{kt}:k=1,\ldots,K\}$ for $t=1$. Most estimated probabilities align closely with the nominal $95$% level, except when $\gamma_{kt}=0$, where coverage is poorer due to the true value residing on the parameter space boundary. Nevertheless, these promising results endorse the XVAE as a fast and robust inference tool for estimating parameters in the max-id process (3) and for Bayesian UQ.
@@ 4 Application to Red Sea surface temperature data
The Red Sea, a biodiversity hot spot, is susceptible to coral bleaching due to climate change and rising SST anomalies (Furby et al., 2013). Corals are unlikely to survive once the temperature exceeds a bleaching threshold annually, which in turn causes disruption in fish migration and slow decline in fish abundance. Here, we analyze and emulate a Red Sea surface temperature dataset, which consists of satellite-derived daily SST estimates at 16,703 locations on a $1/20^{\circ}$ grid from 1985/01/01 to 2015/12/31 (11,315 days in total); see Donlon et al. (2012). This yields about 189 million correlated spatio-temporal data points. Through our analysis, we demonstrate the importance of generating realistic realizations and quantifying uncertainty associated with threshold exceedances—they could for example be used to accurately assess marine heatwave (MHW) risks and identify regions susceptible to coral bleaching by refining the threshold for MHW detection with improved UQ.
Figure 6: In the left two panels, we show empirically-estimated pairwise tail dependence measure $\chi_{0j}(u)$, $u=0.85$, between $\boldsymbol{s}_{0}=(38.104,21.427)$, marked using a red cross, and all $\boldsymbol{s}_{j}\in\mathcal{S}$, from observations and emulated data. In the right two panels, we show the estimated tilting parameters at $K=243$ data-driven knots averaged over time (i.e., $n_{t}^{-1}\sum_{t=1}^{n_{t}}\gamma_{kt}$, $k=1,\ldots,K$), and the estimated tilting parameters averaged over space (i.e., $K^{-1}\sum_{k=1}^{K}\gamma_{kt}$, $t=1,\ldots,n_{t}$) with the best linear regression fit (red line).
Since the daily SST at each location exhibits a clear trend and seasonality across seasons and years and the temporal dependence strength varies at different locations, we first de-trend and remove the seasonality before applying our model (see Section F.1 of the Supplementary Material for details). Despite the successful performance of our XVAE in Section 3 when the data input is not generated from the process (3), we opt to extract monthly maxima from the renormalized data to better comply with the assumed max-infinite divisibility and to enhance the modeling accuracy of marginal distributions for station records. In Section F.2, we perform various goodness-of-fit tests using the generalized extreme value (GEV) distribution and the general non-central $t$ distribution, and find that the former performs better than the latter and it fits the monthly maxima very well at almost all grid points (99.99% of the 16,703 locations). Therefore, we fit a GEV distribution at each location and transform the sitewise records to the Pareto scale on which we then apply the XVAE; see Section F.3 for specifics.
Refer to caption Refer to caption Refer to caption
Figure 7: Observed (left) and emulated (middle) Red Sea SST monthly maxima, for the 1985/01 (top) and 2015/12 (bottom) months. From the emulation maps and QQ plots (right), we see that the emulated fields from the XVAE match the observations very well.
The third panel of Figure 6 displays the locations of the data-driven knots chosen by our algorithm ($K=243$), and the initial radius shared by the Wendland basis functions is 1.2. Figure 7 shows emulated replicates of the original monthly maxima field for the first and last months (1985/01 and 2015/12, respectively). Here, we convert the emulated values back to the original data scale using the estimated GEV parameters fitted from the previous step. Figure 7 demonstrates that the XVAE is able to capture the detailed features of the temperature fields and to accurately characterize spatial dependence, while the QQ-plot shows an almost perfect alignment with the 1-1 line.
@@
Similar to Figure 4, we then estimate $\chi_{h}(u)$ empirically for the original monthly maxima and emulated fields, under the working assumption of stationarity and isotropy. Figure F.2 of the Supplementary Material attests once more that our XVAE characterizes the extremal dependence structure accurately from low to high quantiles. As introduced in Section 2, we also examine the pairwise $\chi$ measure without assuming spatial stationarity. Specifically, we choose the center of the Red Sea $(38.104^{\circ}\mathrm{E}$, $21.427^{\circ}\mathrm{N})$ as a reference point denoted by $\boldsymbol{s}_{0}$. For all observed locations $\boldsymbol{s}_{j}\in\mathcal{S}$ in the Red Sea, we estimate the pairwise $\chi_{0j}(u)$ empirically only using the values from $\boldsymbol{s}_{j}$ and $\boldsymbol{s}_{0}$. The left two panels of Figure 6 include raster plots of the pairwise measure evaluated at the level $u=0.85$, in which the $\chi_{0j}(u)$ values estimated from the observed and emulated data are very similar to each other.
@@
One other major advantage of our XVAE framework is that it allows the dependence parameters $\{\gamma_{k}:k=1,\ldots,K\}$ and $\alpha$ to change over time, which makes the extremal dependence structure non-stationary. The right two panels of Figure 6 show the estimated tilting parameters $\{\gamma_{kt}:k=1,\ldots,K,\;t=1,\ldots,n_{t}\}$ averaged over time/space. We see that the $\gamma_{kt}$ values are generally lower near the coast compared to the interior of the Red Sea, indicating that SST tends to be more heavy-tailed on the coast. Also, $\gamma_{kt}$ tends to increase over time, indicating that extreme events are becoming more localized. This is consistent with the findings in Genevier et al. (2019).
@@
To examine what the SST fields look like without the noise, the left panels of Figure 8 display realizations of the latent field $\{Y_{t}(\boldsymbol{s})\}$ using the fitted XVAE at two time points, 1985/9 and 2015/9. To focus on the extreme values, we transform $Y_{t}(\boldsymbol{s})$ to the original SST scale using the estimated GEV parameters and censor the simulations with a fixed thermal threshold of 31C, resulting in threshold exceedances primarily in the southern region. This is expected: the southern Red Sea experiences higher SSTs compared to the northern area. However, coral reefs in different parts of the Red Sea have developed varying levels of thermal tolerance (Hazra and Huser, 2021). To explore regional variation in marine heatwave (MHW) and coral bleaching risk, we divide the Red Sea into four regions based on Raitsos et al. (2013) and Genevier et al. (2019): North (25.5–30N), North central (22–25.5N), South central (17.5–22N) and South (12.5–17.5N).
A useful quantity is the areal exceedance probability, which represents the spatial extent of a region being simultaneously at extreme risk of MHW. To accurately estimate these joint probabilities (and uncertainties thereof) based on the trained XVAE, we generate 30,000 independent SST emulations for each time point and calculate the total area exceeding the designated threshold of 31C. One could easily define a different, potentially spatially-varying, thermal threshold (e.g., with fixed marginal probability of exceedances). The middle panels of Figure 8 report the results, namely the density of the total area at risk of MHW within each region. The curves for the south central and south regions concentrate around larger areas, while the north central and north regions show little or no exceeded areas, confirming that surface temperatures decrease with increasing latitude. Additionally, the middle panels illustrate that under rising SST conditions, simultaneous exceedances of 31C across larger areas may become more likely over time for all subregions except the north, where 31C still surpasses the highest possible temperature in September 2015.
@@
In order to provide a more detailed description of the spatial extent of areas of joint threshold exceedances at different extreme levels, we then estimate the SST threshold required for an array of fixed spatial extents of exceedances. For each fixed spatial extent, we calculate the minimal threshold needed to reach that area of joint exceedances from each emulated replicate, and we then group all 30,000 estimated thresholds together to derive $95\%$ empirical confidence intervals. We repeat this process for all spatial extents of exceedances in between 100 km2 and $1.4\times 10^{5}$ km2. Note that this can be computed quite fast thanks to the amortized nature of our XVAE. The right panels of Figure 8 illustrate a consistent rise in SST thresholds across all extreme levels from 1985 to 2015. This trend is further confirmed in Figure 9, reporting the same results for a specific area of $5\times 10^{4}$ km2. Slicing the confidence bands at the 31C threshold aligns with the middle panel results, demonstrating that we can provide sensible uncertainty quantification when evaluating the spatial extent of exceedances at any extreme level. This also shows that the joint tail of SSTs exhibits a weakening dependence structure within each subregion, with the spatial extent of exceedances decreasing as the threshold increases and extreme events becoming more localized as they get more extreme. Furthermore, as the spatial extent approaches zero, the threshold estimates represent the highest possible SST for a specific month in a subregion, a valuable metric for studying phytoplankton bloom.
To directly assess the impact of climate change, we determine the threshold necessary for one specific spatial extent of exceedances (i.e., $5\times 10^{4}$ km2) across all September months from 1985 to 2015. Here, we also emulate 30,000 SST fields for each September month in this period. Figure 9 reveals that the fixed-area SST threshold increased steadily by 0.7 in all four subregions on average over the studied time period, corroborating the warming trend in the Red Sea and the localized nature of extremes shown in Figures 6 and 8.

### v4

@@ 4 Simulation study
In this section, we simulate data from five different parametric models that have varying levels of extremal dependence across space. By examining the diagnostics introduced in Section 3.3, we validate the efficacy of our XVAE to analyze and emulate data from both model (3) and misspecified models.
@@ 4.1 General setting
We conduct a simulation study in which data are generated at $n_{s}=2,000$ random locations uniformly sampled over the square $[0,10]\times[0,10]$. We simulate $n_{t}=100$ replicates of the process from each of the following different models:
@@
When simulating from Models IIIV, we first consider time-invariant dependence parameters $\alpha_{t}\equiv 1/2$ and $\boldsymbol{\gamma}_{t}\equiv\boldsymbol{\gamma}$, and attempt to recover the spatial dependence structure; see Figure 2 for the knot locations and $\boldsymbol{\gamma}$ values. Recall that $K$ is the number of basis functions and $\mathcal{D}=\{k:\gamma_{k}=0\}$. We sample the latent variables $\boldsymbol{Z}_{t}$ from the expPS distribution independently for each time replicate. The white noise process $\{\epsilon_{t}(\boldsymbol{s})\}$ follows the same independent Fréchet($0,\tau,1/\alpha_{0}$) distribution with $\tau=1$ and $\alpha_{0}=1/4$.
Figure 2: The left panel presents knot locations used for Models IIIV, and we only show the support of the one Wendland basis function centered at knot in the middle of the domain. Model V uses the same set of knots but the basis functions are not compactly supported. The middle and right panels display the $\gamma_{k}$ values, $k=1,\ldots,K$, used in the expPS variables for Models II and III respectively. The circled knots signify $\gamma_{k}=0$, which induces local AD.
Model I is a stationary and isotropic Gaussian process with a Matérn covariance function. It is known that the joint distribution of the Gaussian process at any two locations $\boldsymbol{s}_{i}$ and $\boldsymbol{s}_{j}$ is light-tailed and leads to AI unless the correlation equals one. For Models IIIV, we simulate data from the max-id model (3) with $K=25$ evenly-spread knots across the grid, denoted by $\{\tilde{\boldsymbol{s}}_{1},\ldots,\tilde{\boldsymbol{s}}_{K}\}$. Setting the range parameter to $r=3$, we use compactly supported Wendland basis functions $\omega_{k}(\boldsymbol{s},r)\propto\{1-d(\boldsymbol{s},\tilde{\boldsymbol{s}}_{k})/r\}^{2}_{+}$ centered at each knot (Wendland, 1995), $k=1,\ldots,K$; see Figure 2. The basis function values are standardized so that for each $\boldsymbol{s}$, $\sum_{k=1}^{K}\omega_{k}(\boldsymbol{s},r)=1$. The main difference between Models II, III and IV lies in the $\gamma_{k}$ values: Model II has no zero $\gamma_{k}$’s (i.e., $|\mathcal{D}|=0$), whereas Model III has a mix of positive and zero $\gamma_{k}$’s, and Model IV has only zero $\gamma_{k}$’s (i.e., $|\mathcal{D}|=K$). By Theorem 3.3, we know Model II gives only local AI and Model IV gives only local AD. In contrast, Model III gives both local AD and local AI. By contrast, Model V adopts the same set of knots but it uses Gaussian radial basis functions which are not compactly supported. Therefore, Model V is the Reich and Shaby (2012) max-stable model.
Models IV gradually exhibit increasingly stronger extremal dependence, and they can help us test whether the XVAE can capture spatially-varying dependence structures that exhibit local AD and/or local AI. Since the proposed process (3) allows $\gamma_{k}$ to change across the different knots ($k=1,\ldots,K$), a well-trained XVAE should be able to differentiate between local AD ($\gamma_{k}=0$) and local AI ($\gamma_{k}>0$).
@@
Additionally, for each space-time simulated dataset, we randomly set aside 100 locations as a validation set. Subsequently, we analyze the dependence structure of the remaining 1,900 locations using both the proposed XVAE (initialized with data-driven knots unless specified otherwise) and a Gaussian process regression with heteroskedastic noise implemented in the R package hetGP (Binois and Gramacy, 2021). We then perform predictions at the 100 holdout locations (see Section B.1 of the Supplementary Material). In the following, we show that both emulators perform well when emulating datasets from Models I and II, but only XVAE appropriately captures heavy tails and AD in Models IIIV.
In Section E.1 of the Supplementary Material, we further examine the XVAE’s ability to capture $\boldsymbol{\gamma}_{t}$ when there is both spatial and temporal nonstationarity. Moreover, in Section E.2 of the Supplementary Material, we simulate data on a regular grid and compare the emulation performance between XVAE and extGAN proposed by Boulaguiem et al. (2022); we see that extGAN has limitations in capturing the extremal dependence appropriately.
@@ 4.2 Emulation results
Figure 3 and Figure E.1 of the Supplementary Material compare emulated replicates from XVAE and hetGP with data replicates from Models IV, while Figure E.2 displays QQ-plots that align well with the 1-1 line in all cases for XVAE but not for hetGP. Since the Gaussian process has much weaker extremal dependence, the resulting $\boldsymbol{\gamma}_{t}$ estimated in (11) after convergence is consistently far greater than $0.1$, indicating light tails in the expPS variables and thus, local AI at all knots. In contrast, Model II exhibits AI across the domain with much smaller $\boldsymbol{\gamma}_{t}$ values (see Figure 2), producing heavier-tailed expPS variables than Model I. As a result, hetGP struggles to capture extremal dependence and the QQ-plot shows its underestimation of large tail values, though Model II still shows only AI.
@@
Figure 4 compares the performance of spatial predictions at the 100 holdout locations. For Model I, hetGP has lower CRPS and MSPE scores, indicating higher predictive power, as expected since the true process is Gaussian. However, the XVAE model still performs quite well in this (misspecified) case. For Models IIV, XVAE uniformly outperforms hetGP. Also, the CRPS and MSPE for hetGP are significantly higher for time replicates with extreme events.
The first three panels of Figure 5 and Figure E.3 of the Supplementary Material compare nonparametric estimates of the upper tail dependence $\chi_{h}(u)$ from the data replicates and emulations at three different distances $h\in\{0.5,2,5\}$ under the working assumption of stationarity. In general, we see that the dependence strength decays as $h$ and $u$ increase, with varying levels of positive limits as $u\rightarrow 1$ for Models IIIV. The results in Figure 5 demonstrate that our XVAE manages to accurately emulate the dependence behavior at both low and high quantiles and the empirical confidence envelopes of $\chi_{h}(u)$ are essentially indistinguishable between the simulated and emulated data.
@@
Figure 6: Initializing the XVAE using the true knots from Model III, we show the estimates of $(\gamma_{1t},\gamma_{3t})^{\rm T}$ (left) and $(\gamma_{5t},\gamma_{12t})^{\rm T}$ (middle) from 1,000 samples generated with the trained decoder $(t=1)$. On the right, we also show the medians, 2.5% and 97.5% quantiles of the $n_{t}$ estimates of $\{\gamma_{kt}:k=1,\ldots,K\}$ for $t=1$, from the decoder (11), in which the 1-1 line is displayed in black for reference.
To showcase the inferential capabilities of our approach, we initialize the XVAE with true knots and rerun it on datasets simulated from Model III. Figure 6 displays $\boldsymbol{\gamma}_{t}$ estimates obtained by running the decoder (i.e., Eq. (11)) $1000$ times at $t=1$. The results highlight the XVAE’s ability to produce accurate estimates of $\boldsymbol{\gamma}_{t}=\{\gamma_{kt}:k=1,\ldots,K\}$, and correctly identify the extremal dependence class, with satisfactory agreement between true and estimated values, accounting for uncertainty. Additionally, we perform a coverage analysis by simulating $99$ more datasets with $n_{s}=2000$ and $n_{t}=100$ from Model III, running the XVAE on each to generate empirical credible intervals for $\boldsymbol{\gamma}_{t}$. Figure E.5 of the Supplementary Material shows the coverage probabilities of $\boldsymbol{\gamma}_{t}$ for $t=1$. Most estimated probabilities align closely with the nominal $95$% level, except when $\gamma_{kt}=0$, where the coverage is poorer due to the true value residing on the parameter space boundary. Nevertheless, these promising results endorse the XVAE as a fast and robust inference tool for estimating parameters in the max-id process (3) and for Bayesian UQ.
@@ 5 Application to Red Sea surface temperature data
The Red Sea, a biodiversity hotspot, is susceptible to coral bleaching due to climate change and rising SST anomalies (Furby et al., 2013). Corals are unlikely to survive once the temperature exceeds a bleaching threshold annually, which in turn causes disruptions in fish migration and slow decline in fish abundance. Here, we analyze and emulate a Red Sea surface temperature dataset, which consists of satellite-derived daily SST estimates at 16,703 locations on a $1/20^{\circ}$ grid from 1985/01/01 to 2015/12/31 (11,315 days in total); see Donlon et al. (2012). This yields about 189 million correlated spatio-temporal data points.
We extract monthly maxima from renormalized data to ensure temporal independence and modeling accuracy of sitewise marginal distributions. Sections F.1F.3 of the Supplementary Material detail how we remove the seasonal trends and how we transform the sitewise records to the Pareto scale on which we then apply the XVAE. The third panel of Figure 7 displays the data-driven knot locations chosen by our algorithm ($K=243$), and the initial radius shared by the Wendland basis functions is 1.2.
@@
Similar to Figure 5, we estimate $\chi_{h}(u)$ empirically for the original monthly maxima and emulated fields, under the working assumption of stationarity and isotropy. Figure F.2 of the Supplementary Material attests once more that our XVAE characterizes the extremal dependence structure accurately from low to high quantiles. Furthermore, we examine the pairwise $\chi$-measures between the center of the Red Sea $(38.104^{\circ}\mathrm{E}$, $21.427^{\circ}\mathrm{N})$ (denoted by $\boldsymbol{s}_{0}$) and all observed locations $\boldsymbol{s}_{j}\in\mathcal{S}$ in the Red Sea. The left two panels of Figure 7 include raster plots of these measures evaluated at the level $u=0.85$, in which the $\chi_{0j}(u)$ values estimated from the observed and emulated data are very similar to each other.
Figure 7: In the left two panels, we show empirically-estimated pairwise tail dependence measure $\chi_{0j}(u)$, $u=0.85$, between $\boldsymbol{s}_{0}=(38.104,21.427)$, marked using a red cross, and all $\boldsymbol{s}_{j}\in\mathcal{S}$, from observations and emulated data. In the right two panels, we show the estimated tilting parameters at $K=243$ data-driven knots averaged over time (i.e., $n_{t}^{-1}\sum_{t=1}^{n_{t}}\hat{\gamma}_{kt}$, $k=1,\ldots,K$), and the estimated tilting parameters averaged over space (i.e., $K^{-1}\sum_{k=1}^{K}\hat{\gamma}_{kt}$, $t=1,\ldots,n_{t}$) with the best linear regression fit (red line).
@@
The right two panels of Figure 7 show estimates of $\{\gamma_{kt}:k=1,\ldots,K,\;t=1,\ldots,n_{t}\}$ averaged over time/space. We see that the $\gamma_{kt}$ values are generally lower near the coast compared to the interior of the Red Sea, indicating SST tends to be more heavy-tailed on the coast. We also observe an upward trend in $\gamma_{kt}$ over time, indicating that extreme events are becoming more localized. This is consistent with the findings in Genevier et al. (2019). It is worth noting that this probabilistic approach to examining the spatio-temporal variation in the dependence structure would not be possible using either hetGP or extGAN.
@@
To focus on extreme values, we transform emulated $\{X_{t}(\boldsymbol{s})\}$ fields to the original SST scale using the fitted marginal distributions and censor the simulations with a fixed thermal threshold of 31C. The left panels of Figure 8 display realizations at two time points, 1985/9 and 2015/9, in which threshold exceedances primarily occur in the southern region. This is expected: the southern Red Sea experiences higher SSTs compared to the northern area. However, coral reefs in different parts of the Red Sea have developed varying levels of thermal tolerance (Hazra and Huser, 2021). To explore regional variation in marine heatwave (MHW) and coral bleaching risk, we divide the Red Sea into four regions based on Raitsos et al. (2013) and Genevier et al. (2019): North (25.5–30N), North Central (22–25.5N), South Central (17.5–22N) and South (12.5–17.5N).
@@
A useful metric is the areal exceedance probability, representing the spatial extent of a region simultaneously at extreme MHW risk. To estimate these joint probabilities and uncertainties, we generate 30,000 independent SST emulations using XVAE for each time point and compute the total area exceeding 31C. Different, potentially spatially-varying thermal thresholds could also be used. The middle panels of Figure 8 show the density of the total area at risk of MHW within each region. The South Central and South regions show larger affected areas, while the North Central and North regions have little to no exceedance, reflecting cooler temperatures with increasing latitude. The results also suggest that under rising SSTs, larger simultaneous exceedances may become more likely over time, except in the North region where 31C remains above the highest possible temperature in 2015.
To further analyze joint exceedances at varying extreme levels, we estimate the SST thresholds required for different fixed spatial extents of exceedances. For each fixed spatial extent, we calculate the minimal threshold needed to reach that area of joint exceedances from each emulated replicate, and we then group all 30,000 estimated thresholds together to derive $95\%$ empirical confidence intervals. We repeat this process for all spatial extents of exceedances in between 100 km2 and $1.4\times 10^{5}$ km2. Note that this can be computed rapidly thanks to the semi-amortized nature of our XVAE. The right panels of Figure 8 reveal a consistent rise in SST thresholds from 1985 (top panel) to 2015 (bottom panel), confirming the warming trend. Slicing the confidence bands at the 31C threshold aligns with the middle panel results. This also shows that within each subregion, the spatial extent of exceedances decreases as the threshold increases and extreme events becomes more localized as they get more extreme. Furthermore, as the spatial extent approaches zero, the threshold estimates represent the highest possible SST for a specific month in a subregion, a valuable metric for studying phytoplankton bloom.
To directly assess the impact of climate change, Figure 9 shows results for specific spatial extent of exceedances (i.e., $5\times 10^{4}$ km2) across all September months from 1985 to 2015. We see that the fixed-area SST threshold has increased steadily by about 0.7 in all four subregions on average over the studied time period, corroborating the warming trend in the Red Sea and the localized nature of extremes shown in Figures 7 and 8.
@@
Through our analysis, we have demonstrated the ability of our XVAE to quantify risk associated with SSTs exceeding critical bleaching thresholds. Our approach not only captures spatial dependence of extreme temperatures but also provides robust UQ. Our findings could support conservation efforts by identifying regions most susceptible to coral bleaching and predicting the potential impact of rising SST on the broader marine environment.